Consider the following Cobb Douglas Production Function:

Y =√K √N

when K = 49 and N = 81

2.2 Is this production function characterised by constant returns to scale? Demonstrate.


1
Expert's answer
2022-04-14T15:31:02-0400

Y = 7 "\cdot" 9 = 63

if we multiply K and N on 2, we will have:

Y = "\sqrt{98 \cdot 162} = 126"

126/63 = 2

So, answer is Yes.